A random sample is drawn from a population of known standard deviation 11.3.


Construct a 90% confidence interval for the population mean based on the information


given (not all of the information given need be used).


a. n = 36, 𝑥̅= 105.2, 𝑠 = 11.2


b. n = 100, 𝑥̅= 105.2, 𝑠 = 11.2



1
Expert's answer
2022-06-01T13:14:37-0400

a.

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:

"CI=(\bar{X}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{X}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(105.2-1.6449\times\dfrac{11.3}{\sqrt{36}}, 105.2+1.6449\times\dfrac{11.3}{\sqrt{36}})"

"=(102.1021, 108.2979)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "102.1021 < \mu < 108.2979," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(102.1021, 108.2979)."


b.

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:

"CI=(\bar{X}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{X}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(105.2-1.6449\times\dfrac{11.3}{\sqrt{100}}, 105.2+1.6449\times\dfrac{11.3}{\sqrt{100}})"

"=(103.3413, 107.0587)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "103.3413 < \mu < 107.0587," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(103.3413, 107.0587)."